Circular No. 8973

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 R3 (LINEAR) An apparently asteroidal object discovered by the LINEAR survey

(and posted on the Minor Planet Center’s ‘NEOCP’ webpage) has been found to show cometary appearance by other astrometric CCD observers. E. Pettarin (Farra d’Isonzo, Italy, 0.40-m f/4.5 reflector, Sept. 7.9 UT) writes that, in a star-rich field, the comet shows a 20′′ tail in p.a. 130^{o}. R. Holmes (Charleston, IL, U.S.A., 0.81-m f/4 astrograph, Sept. 8.1) notes a diffuse, elongated shape with coma diameter 13′′.3. G. Hug (Scranton, KS, U.S.A., 0.56-m reflector, Sept. 8.14–8.15) finds a short, diffuse extension in p.a. ∼ 130^{o}. J. Young (Table Mountain 0.61-m f/16 Cassegrain reflector, Sept. 8.3) reports an 8′′ round, bright coma of mag 17.5 and a poorly defined broad tail of length 16′′ in p.a. 125^{o}–145^{o}. P. Camilleri, G. Sostero, and E. Guido (remotely using a 0.30-m f/5.8 reflector at Trunkey, N.S.W., Australia, Sept. 8.44–8.48) find that nine co-added 120-s unfiltered exposures reveal a coma that is ∼ 12′′ in diameter.

2008 UT  α_{2000}  δ_{2000}  Mag.
Sept.7.12889  18^{h}26^{m}29.^{s}39  +19^{o}10′58′′.2  18.5

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-R51.

T = 2008 Nov. 26.757 TT  ω = 91.^{o}752
Ω = 268.172 }2000.0
q = 1.79822 AU  i = 41.140

NOTICE TO SUBSCRIBERS The Central Bureau (and Minor Planet Center, and ICQ) are now

again able to receive payments for subscriptions, line charges, and donations via credit card (a feature that had unfortunately been disabled for several months). The secure online order form can be accessed at website URL

http://www.cfa.harvard.edu/iau/services/OnLineOrders.html.

V466 ANDROMEDAE Photometry of this cataclysmic variable (cf. IAUC 8971) from S. Shu-

garov and I. Volkov (Astronomical Institute, Slovak Academy of Science, 0.6-m AISAS teelscope; uncertainty 0.03–0.08 mag): Sept. 2.87 UT, U = 12.29, B = 13.16, V = 13.15; 3.08, 12.28, 13.19, 13.09, R_{j} = 13.31.

2008 September 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green